A radio station's channel, such as 100.7 FM or 92.3 FM L, is actually its frequency in megahertz (MHz), where 1 MHz=10^6 Hz and 1Hz=1 s^-1.
Calculate the broadcast wavelength of the radio station 107.9 FM.
Answer #1

frequency of 107.9 FM radiation = 107.9 Mhz

       = 107.9*10^6 Hz

       = 107.9*10^6 s-1

weknow that, c = vl

c= light velocity = 3*10^8 m/s

v = frequency of radiation = 107.9*10^6 s-1

l = wavelength of radiation = x m

3*10^8 = 107.9*10^6*x

x = 2.78 m

Wavelength of radiation (l) = x = 2.78 m